# boxedasideAside

### v1

@@ 3.3 The $p$-rank stratification of the locus of bielliptic curves $\mathcal{B}_{g}$
In both cases, we conclude that $\dim W\leq g-2+f$, which implies $\dim W=g-2+f$, as desired. Consequently, $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension $g-2+f$. ∎
Theorem 1.1 and Corollary 1.2 follow from Theorem 3.12, and Corollaries 3.14 and 3.15.
4 Applications
In this section, we give some explicit construction of families of double covers of an elliptic curve $E$ for which we have control over the $p$-rank. We then look more carefully into the supersingular locus of $\mathcal{B}_{E,g}$ and offer some observations as the consequence of the results we establish in section 3.
4.1 Explicit smooth covers of small $p$-rank
Let $k$ be an algebraically closed field of characteristic $p>2$ and let $E$ be an elliptic curve over $k$ given by the Legrenge form $y^{2}=x(x-1)(x-\lambda)$, for some $\lambda\in k$. Recall that the $p$-rank of a curve $D$ of genus $g$ over $k$ is the semisimple rank of the Frobenius morphism on $H^{1}(D,\mathcal{O}_{D})$, that is the rank over $k$ of the $g^{th}$ semilinear iteration of the Hasse-Witt matrix associated to $D$. In this section, we use the Hasse-Witt matrix to study specific double covers of a given elliptic curve $E$, and construct families of smooth bielliptic curves $D$ for which we have control over its $p$-rank.
We will take the ramification locus of the cover $D\to E$ to consist of $\mathbb{Z}/2\mathbb{Z}$-conjugates, i.e., its ramification points should be of the form
$(x_{1},y_{1}),(x_{1},-y_{1}),\cdots,(x_{n},y_{n}),(x_{n},-y_{n}).$
One way to construct such a cover is to let $D$ be the normalization of $D^{\prime}$, where $D^{\prime}=D_{2}\times_{\mathbb{P}^{1}}E$ is the fibre product of the following curves:
$\begin{split}&E:y^{2}=x(x-1)(x-\lambda),\\ &D_{2}:z^{2}=\prod_{k=1}^{n}(x-x_{k}).\\ \end{split}$ (4)
Let ${\rm Jac}(D)_{\textup{new}}\subset{\rm Jac}(D)$ denote the Prym variety of the cover $D\to E$, such that
${\rm Jac}(D)\sim{\rm Jac}(E)\oplus{\rm Jac}(D)_{\textup{new}}.$ (5)
Equivalently, ${\rm Jac}(D)_{\textup{new}}$ consists of divisors that are negated by the non-trivial element of ${\rm Aut}(D/E)$.
Lemma 4.1.
If $D$ is as above, then:
  1. 1.
  2. 2.
    Let $\iota_{3}=\iota_{1}\circ\iota_{2}$. Consider the order two subgroup groups of ${\rm Aut}(D)$ given by $H_{1}=\{id,\iota_{1}\}$, $H_{2}=\{id,\iota_{2}\}$ and $H_{3}=\{id,\iota_{3}\}$. Then we have the following quotients of $D$:
    1. (a)
      $E=D/H_{1}:y^{2}=x(x-1)(x-\lambda)$;
    2. (b)
      $D_{2}=D/H_{2}:z^{2}=(x-x_{1})\dots(x-x_{n})$, $g_{D_{2}}=\lceil\frac{n}{2}\rceil-1$;
    3. (c)
      $D_{3}=D/H_{3}:w^{2}=x(x-1)(x-\lambda)(x-x_{1})\dots(x-x_{n})$, $g_{D_{3}}=\lceil\frac{n+1}{2}\rceil$. Here $w=yz$.
    In addition, $D_{2}$ and $D_{3}$ are hyperelliptic. The Jacobian of $D$ decomposes, up to isogeny, as ${\rm Jac}(D)\sim{\rm Jac}(E)\oplus{\rm Jac}(D_{2})\oplus{\rm Jac}(D_{3})$, while ${\rm Jac}(D)_{\textup{new}}\sim{\rm Jac}(D_{2})\oplus{\rm Jac}(D_{3})$.
Proof.
These observations follow immediately from the construction in Equation (4), combined with [KR89, Theorem C] and Equation (5). ∎
Now, we use the fibre product to construct double covers of $E$. We first construct a smooth curve $D$ of genus $g=3$ with $f_{D}=0$, which is a double cover of a supersingular curve $E$ in characteristic $p=3$. By Remark 3.9, there are no smooth curves of genus $g=2$ that are double covers of such an $E$.
Lemma 4.2 ($p=3$).
Let $E$ an elliptic curve whose $3$-rank equals 00, i.e., with $f_{E}=0$. Then, there exists a smooth curve $D$ of genus $3$ with $f_{D}=0$.
Proof.
We use the fact that in characteristic $3$, $y^{2}=x^{4}+ax^{3}+bx^{2}+cx+d$ is a supersingular elliptic curve if and only if $b=0$ and $x^{4}+ax^{3}+bx^{2}+cx+d=0$ has $4$ distinct roots. This is because we need $4$ distinct roots to ensure that the curve is smooth and then the Hasse-invariant is $b$.
We write $i\in\mathbb{F}_{9}$ for any of the roots of $x^{2}=-1$. Take
$E:y^{2}=x(x-i-2)(x-2i-2)(x-1)=x^{4}+x^{3}+x$
Notice that over an algebraically closed field of characteristic $3$, there is up to isomorphism a unique supersingular elliptic curve, so without lost of generality, we can assume $E$ is given by the above equation. Let $D$ be the normalization of $D_{2}\times_{\mathbb{P}^{1}}E$, where
$D_{2}:z^{2}$ $=x(x-i-2)(x-2i)(x-i-1)=x^{4}+2ix^{3}+ix.$
By lemma 4.1, We see that ${\rm Jac}(D)\sim{\rm Jac}(E)\oplus{\rm Jac}(D_{2})\oplus{\rm Jac}(D_{3})$, where
$D_{2}:$ $z^{2}=x(x-i-2)(x-2i)(x-i-1)$ $=x^{4}+2ix^{3}+ix,$
$D_{3}:$ $w^{2}=(x-2i-2)(x-1)(x-2i)(x-i-1)$ $=x^{4}+(i+2)x^{3}+(2i+2)x+1.$
We see that $E$, $D_{2}$ and $D_{3}$ are supersingular and $g_{D}=3$. Thus $D$ is a genus $3$ double cover of $E$ and the $3$-rank of $D$ is 00. ∎
Next, for a general $p>2$, we want to explicitly construct a one-parameter family of $D$ whose $p$-rank is small. We first construct families of $D_{2}$ with extra symmetry. The authors of [DH17] considered the curve $z^{2}=x^{n}+1$ under the condition that $n$ and $\frac{n-1}{2}$ are both prime. Further, assuming $p\not\equiv 0,1\pmod{n}$, they show that the curve has $p$-rank zero. We look at the same curve $z^{2}=x^{n}+1$, without requiring $n$ or $\frac{n-1}{2}$ to be prime.
Lemma 4.3.
Let $t\in k$ be such that $t\neq 0$. Consider the curve $D$ given by $z^{2}=x^{n}-t^{n}$. If $p\nmid n$ and $p\not\equiv 1\pmod{n}$, then $D$ is not ordinary. Moreover, if $p\equiv-1\pmod{n}$, then $D$ is superspecial, so in particular it has $p$-rank 00.
Proof.
Let $g=g(D)=\lceil\frac{n}{2}\rceil-1$ and $M$ be the Hasse-Witt matrix of $D$. Note that $D$ is ordinary if and only if the $p$-rank of $D$ is $g$, which is if and only if the rank of $M$ is $g$.
It is known that $H^{1}(D,\mathcal{O}_{D})$ has basis given by $\frac{z}{x^{k}}$ for $1\leq k\leq g$. Notice that ${\rm Aut}(D)$ contains $\mathbb{Z}/n\mathbb{Z}$, the action given by $(x,z)\to(\zeta_{n}x,z)$. Since $p\nmid n$, this action leads to the decomposition
$H^{1}(D,\mathcal{O}_{D})=\bigoplus_{i=0}^{n-1}H^{1}(D,\mathcal{O}_{D})_{i}.$
Notice that for $n-g\leq i\leq n-1$, we have $H^{1}(D,\mathcal{O}_{D})_{i}=\langle\frac{z}{x^{n-i}}\rangle$ and for $10\leq i\leq n-g-$ 1, we have $H^{1}(D,\mathcal{O}_{D})_{i}=0$. Moreover, $M$ maps $H^{1}(D,\mathcal{O}_{D})_{i}$ to $H^{1}(D,\mathcal{O}_{D})_{pi\pmod{n}}$. This means the rank of $M$ equals #$\#S$, where $S\subseteq\{1,2,\dots,g\}$ is the following set
$S$ =$=\{n-g\leq i\leq n-1\mid n-g\leq pi\pmod{n}\leq n-1\}$
=$=\{1\leq i\leq g\mid 1\leq pi\pmod{n}\leq g\}.$
Let $a\equiv p\pmod{n}$, so $12\leq a\leq n-$ 1.
  1. 1.
    If $g+1\leq a\leq n-1$. Then $1\notin S$, so $D$ is not ordinary.
  2. 2.
    If $2\leq a\leq g$. Let $i=\lceil\frac{g+1}{a}\rceil$, so $1\leq i\leq g$. Moreover, $g+1\leq ai<g+1+a\leq 2g+1\leq n$. This means that $i\notin S$, so $D$ is not ordinary.
In both cases, $D$ is not ordinary. In addition, if $p\equiv-1\pmod{n}$, then $S=\emptyset$. Therefore, $M$ has rank 00 and hence the $p$-rank of ${\rm Jac}(D)$ is also 00. ∎
Now, we are using $D$ as $D_{2}$ and construct double covers of $E$ with small $p$-rank.
Proposition 4.4.
Let $k$ be an algebraically closed field of characteristic $\text{char}(k)=p$ and $t\in k^{*}$. For any $\lambda\in k$, let $E:y^{2}=x(x-1)(x-\lambda)$ be an elliptic curve over $k$, $D_{2}:z^{2}=f(x)$ be a (hyper)elliptic curve over $k$, and $D$ be the normalization of $E\times_{\mathbb{P}^{1}}D_{2}$.
  1. 1.
    Suppose $p\not\equiv 1\pmod{n}$ and $p\nmid n$. If $D_{2}$ is given by $z^{2}=x^{n}-t^{n}$, then ${\rm Jac}(D)_{\rm new}$ is not ordinary. Moreover, if $p\equiv-1\pmod{n}$, then $f_{D}\leq\lceil\frac{n+1}{2}\rceil$.
  2. 2.
    Suppose $n$ is odd, and $p\equiv-1\pmod{n+3}$ and $p\nmid n+3$. Let $\zeta_{n+3}$ be a primitive $(n+3)^{th}$ root of unity. Let $x_{1},\cdots,x_{n}\in\mathbb{P}^{1}(k)$ be determined by the following condition: Let $L$ be the unique linear fractional transformation that sends $(t,\zeta_{n+3}t,\zeta_{n+3}^{2}t)$ to $(0,1,\lambda)$ and $x_{i}=L(\zeta_{n+3}^{i+2}t)$. If $D_{2}:z^{2}=(x-x_{1})\cdots(x-x_{n})$, then $f_{D}\leq\frac{n-1}{2}$.
Proof.
  1. 1.
    We have that $f_{{\rm Jac}(D)_{\textup{new}}}=f_{D_{2}}+f_{D_{3}}$, where $D_{3}:w^{2}=x(x-1)(x-\lambda)(x^{n}-t^{n})$. In particular, $f_{D_{3}}\leq g_{D_{3}}\leq\lceil\frac{n+1}{2}\rceil$. By Lemma 4.3, we have that $f_{D_{2}}\leq g_{D_{2}}-1$, hence ${\rm Jac}(D)_{\textup{new}}$ is not ordinary. Moreover, when $p\equiv-1\pmod{n}$, then $f_{D_{2}}=0$, so $f_{{\rm Jac}(D)_{\textup{new}}}=f_{D_{3}}\leq g_{D_{3}}=\lceil\frac{n+1}{2}\rceil$.
  2. 2.
    We have that $f_{{\rm Jac}(D)_{\textup{new}}}=f_{D_{2}}+f_{D_{3}}$ where $D_{3}:w^{2}=x(x-1)(x-\lambda)(x-x_{1})\cdots(x-x_{n})$. Since $n$ is odd, the branching points of $D_{3}$ are exactly $0,1,\lambda,x_{1},\cdots,x_{n}$ (it is not branched at $\infty$). By the change of variable $x^{\prime}=L^{-1}x$, we see that $D_{3}$ is isomorphic to the curve given by $w^{2}=x^{\prime n+3}-t^{n+3}$. By Lemma 4.3, we see that when $p\equiv-1\pmod{n+3}$, then $f_{D_{3}}=0$. Hence, $f_{{\rm Jac}(D)_{\textup{new}}}\leq f_{D_{2}}\leq\frac{n-1}{2}$.
Proposition 4.5.
Suppose we are given an elliptic curve $E:y^{2}=x(x-1)(x-\lambda)$ and $n\geq 2$ such that $p\not\equiv 1,n-2\pmod{2(n-1)}$ and $p\nmid 2(n-1)$. Let $D$ be the normalization of the fibre product of $y^{2}=x(x-1)(x-\lambda)$ and $z^{2}=x(x^{n-1}-1)$. Then $g(D)=n$ and ${\rm Jac}(D)_{\rm new}$ is not ordinary. In addition, if $p\equiv-1\pmod{2(n-1)}$, then ${\rm Jac}(D)_{\rm new}$ has p-rank at most $\lceil\frac{n-1}{2}\rceil-1$.
Proof.
In this case we have $D_{1}:z^{2}=x(x^{n-1}-1)$ and $D_{2}:w^{2}=(x-\lambda)(x^{n-2}+x^{n-3}+\dots+1)$. So $g_{1}=g(D_{1})=\lceil\frac{n}{2}\rceil-1$, $g_{2}=g(D_{2})=\lceil\frac{n-1}{2}\rceil-1$ and $g=1+g_{1}+g_{2}=n$. If the we show that $D_{1}$ is not ordinary, then ${\rm Jac}_{\rm new}(D)$ will also be non-ordinary. Let $M$ be the Hasse Witt matrix of $D_{1}$.
It is known that $H^{1}(D_{1},\mathcal{O}_{D_{1}})$ has basis given by $\frac{z}{x^{k}}$ for $1\leq k\leq g_{1}$. Notice that ${\rm Aut}(D_{1})$ contains $\mathbb{Z}/2(n-1)\mathbb{Z}$, the action given by $(x,z)\to(\zeta_{n-1}x,\zeta_{2(n-1)}z)$. Since $p\nmid 2(n-1)$, this action leads to the decomposition
$H^{1}(D_{1},\mathcal{O}_{D_{1}})=\bigoplus_{i=0}^{2n-3}H^{1}(D_{1},\mathcal{O}_{D_{1}})_{i}.$
Notice that for odd $i$ in the range $32n-2g_{1}-1\leq i\leq 2n-$ 3, we have $H^{1}(D_{1},\mathcal{O}_{D_{1}})_{i}=\langle\frac{z}{x^{(2n-1-i)/2}}\rangle$. For other $i$, $30\leq i\leq 2n-$ 3 we have $H^{1}(D_{1},\mathcal{O}_{D_{1}})_{i}=0$. $\phi$ maps $H^{1}(D_{1},\mathcal{O}_{D_{1}})_{i}\to H^{1}(D_{1},\mathcal{O}_{D_{1}})_{pi\pmod{2(n-1)}}$. Hence, the rank of $M$ is given by
#$\#\{i:2n-2g_{1}-1\leq i\leq 2n-3,i\text{ odd},\;2n-2g_{1}-1\leq pi\mod 2(n-1)\leq 2n-3\}$
=#$=\#\{i:\;1\leq i\leq 2g_{1}-1,i\text{ odd},\;1\leq pi\mod 2(n-1)\leq 2g_{1}-1\}.$
Denote this set as $S$. Clearly $S\subseteq\{1,3,\dots,2g_{1}-1\}$. Let $a\equiv p\mod 2(n-1)$, so $a$ is odd and $33\leq a\leq 2n-$ 3 and $a\neq n-2$.
  • $\cdot$
    Case 1: $32g_{1}+1\leq a\leq 2n-$ 3. Then $1\notin S$. So #$\#S\leq g_{1}-1$.
  • $\cdot$
    Case 2: $3\leq a\leq\frac{n}{2}$. Let $i=\lceil\frac{2g_{1}}{a}\rceil$, so $21\leq i\leq 2g_{1}-$ 2. Moreover, $12g_{1}\leq ai<a(i+1)<2g_{1}+2a\leq 2g_{1}+n\leq 2n-$ 1. Thus $i,i+1\notin S$, so #$\#S\leq g_{1}-1$.
  • $\cdot$
    Case 3: $\frac{n}{2}<a\leq 2g_{1}-1$ and $a\neq n-2$, so $n<2a\leq 4g_{1}-2\leq 2n-4$. This forces $n+1\leq 2a\leq 2n-6$, that is, $34\leq(2n-2)-2a\leq n-$ 3. Let $a_{1}=2n-2-2a$. Assume for the sake of contradiction that $S=\{1,3,\dots,2g_{1}-1\}$. Say $b=(2g_{1}-1)a\pmod{2n-2}$. Since $2g_{1}-1\in S$, we know that $b\in S$. Let $j=\lceil\frac{2g_{1}-b}{a_{1}}\rceil$ and $i=2g_{1}-1-2j$. We have $1\leq j<1+\frac{2g_{1}-1}{4}$ and hence $23\leq i\leq g_{1}-$ 2 and $i$ is odd, that is $i\in S$. Next, note that
    $ai=a(2g_{1}-1-2j)\equiv b+a_{1}j\pmod{2n-2}.$
    Finally we have $\frac{2g_{1}-b}{a_{1}}\leq j<\frac{2g_{1}-b}{a_{1}}+1$, i.e., $22g_{1}\leq ja_{1}+b<2g_{1}+a_{1}\leq(n-1)+(n-3)<2n-$ 2. This contradicts the fact that $i\in S$. Therefore $S\neq\{1,3,\dots,2g_{1}-1\}$ and hence #$\#S\leq g_{1}-1$.
Since #$\#S<g_{1}$, we see that $D_{1}$ is not ordinary. In addition, if $p\equiv-1\pmod{n}$, then $S=\emptyset$. Therefore, $M$ has rank 00 and hence the $p$-rank of $D_{1}$ is also 00. It follows that the $p$-rank of ${\rm Jac}_{\rm new}(D)$ is equal to the $p$-rank of $D_{2}$, which is at most $g_{2}=\lceil\frac{n-1}{2}\rceil-1$. ∎
4.2 Supersingular double covers of $E$
Let $k$ be an algebraically closed field of characteristic $p>2$, as above. Recall that a smooth curve $D$ of genus $g$ over $k$ is called supersingular if its Jacobian ${\rm Jac}(D)$ is isogenous to a product $E^{g}$, where $E$ is one of the finitely many supersingular elliptic curves over $k$. Equivalently, its Newton polygon, the combinatorial way to encode the isogeny classes of ${\rm Jac}(D)[p^{\infty}]$, is of the form $\left({\frac{1}{2},\frac{1}{2},\ldots,\frac{1}{2}}\right)$. In particular, if $D$ is supersingular, it has $p$-rank 00. We denote by $\mathcal{M}_{g}^{ss}\subseteq V_{0}(\mathcal{M}_{g})$ the locus of all supersingular smooth curves of genus $g$ in characteristic $p$.
@@ 4.2 Supersingular double covers of $E$
    There are only finitely many supersingular curves of genus $2$ that are double covers of an elliptic curve (necessarily a supersingular one). This follows from the fact that $V_{0}(\mathcal{B}_{E,2})$ is a 00-dimensional locus and the fact that for $g=2$, being supersingular is equivalent to having $p$-rank 00. See Remark 3.7 for more details.
@@
    The locus of supersingular curves of genus $3$ that are double covers of an elliptic curve is $1$-dimensional. Indeed, we have shown that $V_{0}(\mathcal{B}_{E,3})$ is a $1$-dimensional locus consisting entirely of supersingular genus 3 curves. Specifically, the classification of eligible Newton polygons of height $62g=$ 6 indicates that any genus $3$ curve with $p$-rank 00 and a map to an elliptic curve is automatically supersingular. See also [Oor91, Theorem 1.12, Remark 3.3].
@@
    The generic point of any component of the supersingular locus $\mathcal{M}_{4}^{ss}$ of $\mathcal{M}_{4}$ does not correspond to a smooth curve that is a double cover of an elliptic curve. Kudo-Harashita-Senda in [KHS20] or Pries in [Pri24] show the existence of a smooth curve of genus $4$ which is supersingular in any characteristic $p>0$, so that $\mathcal{M}_{4}^{ss}$ is non-empty; any of its components is either $3$-dimensional or $4$-dimensional. On the other hand, we have shown that for any supersingular elliptic curve $E$, $V_{0}(\mathcal{B}_{E,4})$ is pure of dimension $2$, which implies the result.
    Note that there are only $2$ possible Newton polygons for a curve $D$ of genus $4$ with $p$-rank 00 which is a double cover of an elliptic curve. Namely, the Newton polygon of $D$ is either $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ or the supersingular one $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2})$.
    Moreover, for $p>3$, the supersingular curve that Kudo-Harashita-Senda found is a Howe curve $H$, i.e., a normalization of the fiber product of two elliptic curves $E_{i},i=1,2$, such that the degree $2$ maps $E_{i}\to\mathbb{P}^{1}$ share exactly one branch point. In particular, $H$ is a double cover of $E_{1}$ and $E_{2}$. If we take now $E=E_{i}$ for any $i=1,2$, this shows that the locus $\mathcal{B}_{E,4}^{ss}$ of supersingular curves in $\mathcal{B}_{E,4}$ has dimension
    $1\leq\dim\mathcal{B}_{E,4}^{ss}\leq 2,$
    because of the purity theorem [DJO00, Theorem 4.1].
@@
When $g=5$, we consider the fiber product of an elliptic curve and an hyperelliptic curve of $g=2$. We fix $p\equiv 11\pmod{12}$.
@@
By Lemma 4.3, we see that when $p\equiv 11\pmod{12},E^{\prime},D^{\prime}$ both have $p$-rank 00. Let $D^{\prime\prime}$ be the normalization of the fiber product of $E^{\prime},D^{\prime}$ over $\mathbb{P}^{1}$. Then for $p\neq 2$, the $p$-rank of $D^{\prime\prime}$ is the $p$-rank of $w^{2}=(x^{2}+1)(x^{4}+x^{2}+1)$. Following the strategy of [KHS20], we parameterize curves isomorphic to $E^{\prime},D^{\prime}$ and solve for the parameters that give fiber product $p$-rank zero.
@@
Notice that $E_{u}\cong E^{\prime}$, $x=\frac{x_{1}+u}{ux_{1}+1},y=\frac{y_{1}}{(ux_{1}+1)^{2}}$. $D_{v}\cong D^{\prime}$, with $x=\frac{x_{1}+v}{vx_{1}+1},y=\frac{y_{1}}{(vx_{1}+1)^{3}}$. Then let $f(x)=(x+u)^{2}+(ux+1)^{2})((x+v)^{4}+(x+v)^{2}+1)$. The Hasse-Witt matrix of $D_{u,v}$ is a $2\times 2$ matrix $M$ with $M_{i,j}=c_{pj-i}$, where $c_{pj-i}$ is the coefficient of $x^{pj-i}$ in $f(x)^{\frac{p-1}{2}}$. Then, $D_{u,v}$ is supersingular if and only if $MM^{(p)}=0$. By [KHS20, proposition 2.2], this is equivalent to
@@
[…]
Since $a,b,c,d$ are polynomials in terms of $u,v$, we plug in $u,v$ into the expression of $a,b,c,d$. Then, the three equations in Equation 6 becomes $h_{i}(u,v)$ for $1\leq i\leq 3$.
Therefore, the $(u,v)$ such that $E_{u}\times_{\mathbb{P}^{1}}D_{v}$ is supersingular are the solutions of $h_{i}(u,v)=0,1\leq i\leq 3$. Also we need to exclude $u=\pm 1,v=\pm 1$ and the $(u,v)$ such that $((x+u)^{2}+(ux+1)^{2})$ and $((x+v)^{4}+(x+v)^{2}+1)$ have a common root. We can find a root for $(u,v)$ in $\mathbb{F}_{p}$ for $p<1000$, except $p=107,443,491$. So if we assume $p\neq 107,443,491,p\leq 1000,p\equiv 11\pmod{12}$, we can construct $D=E_{u}\times_{\mathbb{P}^{1}}D_{v}$ such that $f_{D}=0$. By the above analysis, this implies that $D$ is a double cover of $E_{u}$ that is supersingular.
@@
    Similarly to the cases $(g=3)$ and $(g=4)$, we can see that there are only $3$ possible Newton polygons for a curve $D$ of genus $5$ with $p$-rank 00 which is a double cover of an elliptic curve. Namely, the Newton polygon of $D$ is one of the following: $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{2},\frac{1}{2},\frac{3}{4},\frac{3}{4},\frac{3}{4},\frac{3}{4})$, $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{2}{3},\frac{2}{3},\frac{1}{3})$, or $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2})$, the supersingular one.
    Therefore, for any $E$ such that there is a supersingular double cover $D\to E$ of genus $5$, it follows that
    $1\leq\dim\mathcal{B}_{E,5}^{ss}\leq 3$
@@
    In particular, for $E$ as in Example 4.6, we find that there is at least a $1$-dimensional family of smooth curves $D$, which are double covers of $E$.

### v2

@@ 3.3 The $p$-rank stratification of the locus of bielliptic curves $\mathcal{B}_{g}$
In both cases, we conclude that $\dim W\leq g-2+f$, which implies $\dim W=g-2+f$, as desired. Consequently, $V_{f}(\overline{\mathcal{B}}_{g})$ is pure of dimension g2+fg-2+f. ∎
Theorem 1.1 and Corollary 1.2 follow from Theorem 3.12, and Corollaries 3.14 and 3.16.
4 Supersingular double covers of $E$
In this section, we look more carefully into the supersingular locus of $\mathcal{B}_{E,g}$ and offer some observations as a consequence of the results we establish in Section 3. Let $k$ be an algebraically closed field of characteristic $p>2$, as above. Recall that a smooth curve $D$ of genus $g$ over $k$ is called supersingular if its Jacobian ${\rm Jac}(D)$ is isogenous to a product $E^{g}$, where $E$ is one of the finitely many (up to isomorphism) supersingular elliptic curves over $k$. Equivalently, its Newton polygon, the combinatorial way to encode the isogeny classes of ${\rm Jac}(D)[p^{\infty}]$, is of the form $\left({\frac{1}{2},\frac{1}{2},\ldots,\frac{1}{2}}\right)$. In particular, if DD is supersingular, it has $p$-rank 0. We denote by $\mathcal{M}_{g}^{\rm ss}\subseteq V_{0}(\mathcal{M}_{g})$ the locus of all supersingular smooth curves of genus $g$ in characteristic $p$.
@@ 4 Supersingular double covers of $E$
    There are only finitely many supersingular curves of genus $2$ that are double covers of an elliptic curve (necessarily a supersingular one). This follows from the fact that $V_{0}(\mathcal{B}_{E,2})$ is a 0-dimensional locus and the fact that for $g=2$, being supersingular is equivalent to having $p$-rank 0. See Remark 3.7 for more details.
@@
    The locus of supersingular curves of genus $3$ that are double covers of an elliptic curve is $1$-dimensional. Indeed, we have shown that $V_{0}(\mathcal{B}_{E,3})$ is a $1$-dimensional locus consisting entirely of supersingular genus 3 curves. Specifically, the classification of eligible Newton polygons of height $2g=6$ indicates that any genus-$3$ curve with $p$-rank 0 and a map to an elliptic curve is automatically supersingular. See also [Oor91, Theorem 1.12, Remark 3.3].
@@
    The generic point of any component of the supersingular locus $\mathcal{M}_{4}^{\rm ss}$ of $\mathcal{M}_{4}$ does not correspond to a smooth curve that is a double cover of an elliptic curve. Kudo-Harashita-Senda in [KHS20] or Pries in [Pri24] show the existence of a smooth curve of genus $4$ which is supersingular in any characteristic $p>0$, so that $\mathcal{M}_{4}^{\rm ss}$ is non-empty; any of its components is either $3$-dimensional or $4$-dimensional. On the other hand, we have shown that for any supersingular elliptic curve $E$, $V_{0}(\mathcal{B}_{E,4})$ is pure of dimension $2$, which implies the result.
    Note that there are only $2$ possible Newton polygons for a curve $D$ of genus $4$ with $p$-rank 0 which is a double cover of an elliptic curve. Namely, the Newton polygon of DD is either $\left(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{2}{3},\frac{2}{3},\frac{2}{3}\right)$ or the supersingular one $\left(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2}\right)$.
    Moreover, for $p>3$, the supersingular curve that Kudo-Harashita-Senda found is a Howe curve $H$, i.e., a normalization of the fiber product of two elliptic curves $E_{i},i=1,2$, such that the degree $2$ maps $E_{i}\to\mathbb{P}^{1}$ share exactly one branch point. In particular, $H$ is a double cover of $E_{1}$ and $E_{2}$. If we take now $E=E_{i}$ for any $i=1,2$, this shows that the locus $\mathcal{B}_{E,4}^{\rm ss}$ of supersingular curves in $\mathcal{B}_{E,4}$ has dimension 1≤dimB_E, 4^ss≤2, because of the purity theorem [DJO00, Theorem 4.1].
@@
When $g=5$, we consider the fiber product of an elliptic curve and a hyperelliptic curve of $g=2$. We fix $p\equiv 11\pmod{12}$.
@@
By Lemma 5.2, we see that when $p\equiv 11\pmod{12},E^{\prime},D^{\prime}$ both have $p$-rank 0. Let $D^{\prime\prime}$ be the normalization of the fiber product of $E^{\prime},D^{\prime}$ over $\mathbb{P}^{1}$. Then the $p$-rank of $D^{\prime\prime}$ is the $p$-rank of $w^{2}=(x^{2}+1)(x^{4}+x^{2}+1)$. Following the strategy of [KHS20], we parametrize curves isomorphic to $E^{\prime},D^{\prime}$ and solve for the parameters that give fiber product $p$-rank zero.
@@
Note that $E_{u}\cong E^{\prime}$, $x=\frac{x_{1}+u}{ux_{1}+1},y=\frac{y_{1}}{(ux_{1}+1)^{2}}$. Observe also $D_{v}\cong D^{\prime}$, with $x=\frac{x_{1}+v}{vx_{1}+1},y=\frac{y_{1}}{(vx_{1}+1)^{3}}$. Then let $f(x)=(x+u)^{2}+(ux+1)^{2})((x+v)^{4}+(x+v)^{2}+1)$. The Hasse-Witt matrix of $D_{u,v}$ is a $2\times 2$ matrix $M$ with $M_{i,j}=c_{pj-i}$, where $c_{pj-i}$ is the coefficient of $x^{pj-i}$ in $f(x)^{\frac{p-1}{2}}$. Then, $D_{u,v}$ is supersingular if and only if $MM^{(p)}=0$. By [KHS20, proposition 2.2], this is equivalent to
@@
$a=c_{p-1},\hskip 14.22636ptb=c_{2p-1},\hskip 14.22636ptc=c_{p-2},\hskip 14.22636ptd=c_{2p-2}$
Since $a,b,c,d$ are polynomials in terms of $u,v$, we plug in $u,v$ into the expression of $a,b,c,d$. Then, the three equations in (6) become $h_{i}(u,v)$ for $1\leq i\leq 3$.
Therefore, the $(u,v)$ such that $E_{u}\times_{\mathbb{P}^{1}}D_{v}$ is supersingular are the solutions of $h_{i}(u,v)=0$ for $1\leq i\leq 3$. Also we need to exclude $u=\pm 1,v=\pm 1$ and the $(u,v)$ such that $((x+u)^{2}+(ux+1)^{2})$ and $((x+v)^{4}+(x+v)^{2}+1)$ have a common root. We can find a root for $(u,v)$ in $\mathbb{F}_{p}$ for $p<1000$, except $p=107,443,491$. So if we assume $p\neq 107,443,491,p\leq 1000,p\equiv 11\pmod{12}$, we can construct $D=E_{u}\times_{\mathbb{P}^{1}}D_{v}$ such that $f_{D}=0$. By the above analysis, this implies that DD is a double cover of $E_{u}$ that is supersingular.
@@
    Similarly to the cases $(g=3)$ and $(g=4)$, we can see that there are only $3$ possible Newton polygons for a curve $D$ of genus $5$ with $p$-rank 0 which is a double cover of an elliptic curve. Namely, the Newton polygon of DD is one of the following: $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{2},\frac{1}{2},\frac{3}{4},\frac{3}{4},\frac{3}{4},\frac{3}{4})$, $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{2}{3},\frac{2}{3},\frac{1}{3})$, or $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2})$, the supersingular one.
@@
    Therefore, for any $E$ such that there is a supersingular double cover $D\to E$ of genus $5$, it follows that 1≤dimB_E, 5^ss ≤3. In particular, for $E$ as in Example 4.1, we find that there is at least a $1$-dimensional family of smooth curves $D$, which are double covers of $E$.